The larger of two numbers is 5 more than twice the smaller. If the smaller is subtracted from the larger, the result is 12. Find the two numbers?
Answer by JulietG(1812)

A = 2B + 5
A - B = 12
Easier to do if the terms are on the same side for both equations, so we'll add B to each side in the second equation.
A = B + 12
A = 2B + 5
Since both equations are equal to A, they are equal to each other
B + 12 = 2B + 5
Subtract B from each side
12 = b + 5
Subtract 5 from each side
7 = B
.
If B is 7, then A is (2*7)+5
A is 14+5 = 19